OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2 - 6x + 8 = y A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3) A, B, C, or D? Please explain :)

OpenStudy (anonymous):

divide the number in front of x by 2 and then square it, then add and subtract that number into the equation (-6/2)^2=9 x^2-6x+9-9+8 factor first three terms and add -9+8 y=(x-3)^2-1 this is vertex form y=a(x-h)^2+k where the vertex is (h,k) so the vertex is (3,-1)

OpenStudy (anonymous):

Thanks a lot :)